### v2

@@ 9 Conclusion
In this study, we evaluated the performance and environmental impact of three pre-trained language models: LLaMA 3-8B, T5, and BART. While LLaMA 3-8B exhibited lower ROUGE and METEOR scores in comparison to the other two models, it demonstrated comparable semantic similarity performance. However, the environmental cost of LLaMA 3-8B is considerably higher, and therefore, one should try to use the smaller models unless absolutely necessary. These findings highlight the trade-offs between model performance and environmental sustainability, emphasizing the need for more energy-efficient approaches in the development and deployment of large-scale language models.

### v3

@@ 9 Conclusion
We have evaluated the performance and environmental impact of three pre-trained language models: LLaMA 3-8B, T5, and BART. Although LLaMA 3-8B produced lower ROUGE and METEOR scores than the other two models, it achieved comparable performance in terms of semantic similarity. However, its environmental cost was significantly higher. This shows that smaller fine-tuned models remain effective for summarization tasks. Future research should prioritize reducing the carbon footprint of language models.
9.0.1 Acknowledgements.
This work is supported by the project CRG/2021/000803, funded by the Department of Science and Technology, Government of India.